In 2004 Jeremiah is five times the age of his cousin Karen. In 2009 he is three times her age. How old was Jeremiah in 2004?
Answer by lwsshak3(11628)

In 2004 Jeremiah is five times the age of his cousin Karen. In 2009 he is three times her age. How old was Jeremiah in 2004?
**
In2004
let x=Karen's age
5x=Jeremiah's age
..
In2009 (5 yrs hence)
x+5=Karen's age
5x+5=Jeremiah's age
Jeremiah is 3 times the age of Karen
5x+5=3(x+5)
5x+5=3x+15
2x=10
x=5
5x=25
Jeremiah's age in 2004=25